# Learning to Generalize towards Unseen Domains via a Content-Aware Style Invariant Model for Disease Detection from Chest X-rays

### 04 Jan 2024 v3

@@ I Introduction
In this work, we propose a novel domain-agnostic, content-biased, and style-invariant DG framework that can be trained in single or multi-source domain settings without explicit domain labels. We employ style randomization modules (SRMs) at the image and feature levels to hierarchically augment the style while preserving the pathology-specific content traits. Utilizing the prior maximum and minimum pixel value knowledge of a normalized CXR image, we construct a set for sampling mean and standard deviation to use them for switching style characteristics at the image level. For feature-level style augmentation, we deploy a learnable SRM module that generates the pixel-wise style embeddings instead of setting pre-defined parameters, i.e., channel-wise mean and standard deviations as affine parameter values [27], for style manipulation. In addition, we employ two consistency regularization losses to enforce the consistency of semantic global feature maps and the predicted probability distributions, respectively, between the original CXR and the style-perturbed CXR. The variability of styles between different perturbations of the same CXR with content information intact makes the style attribute unreliable and makes the model focus more on pathology-specific content in making a decision. The contributions of this paper are summarized as follows:
@@ II Related work
The feature-level randomization includes the perturbation of feature statistics from intermediate convolutional layers of the backbone model. Nuriel et al.[29] applied adaptive instance normalization to swap the style statistics between activations of permuted and non-permuted samples in a mini-batch to reduce the representation of global statistics in the classifier model. However, they merely swap the style statistics, limiting the sample diversity to the available source domains. Zhou et al.[19] proposed a novel parameter-free module MixStyle that synthesizes new domains by probabilistically mixing feature statistics between instances. Nevertheless, insufficient diversity still exists due to reliance on source domains. Nam et al.[21] utilized adaptive instance normalization along with adversarial learning to reduce the intrinsic style bias. However, the style-biased network increases training cost and time. Tang et al.[30] proposed two novel modules named CrossNorm and SelfNorm for swapping feature statistics and recalibrating the statistics through attention to both diversify and bridge the gap between source and target domains. Though the SelfNorm is dynamic, the CrossNorm is limited to training data. Wang et al.[22] developed an encoder-decoder-like architecture named FSR to integrate random noise with the original style parameters to generate new style embeddings. However, they use the pre-defined parameters, i.e., channel-wise mean and standard deviation, as the style embeddings. Instead of pre-defined parameters, pixel-wise learnable style parameters may yield more diverse and hard style statistics, resulting in a more robust DG model.
@@ III-B1 Instance Normalization
Figure 2: Overview of the proposed framework. The style statistics of the input CXR are randomized with the randomly sampled mean and standard deviation from the set of values constructed from the prior knowledge of possible maximum and minimum values. Both CXRs, with or without style randomized, are passed to the shared feature extractor, DenseIBN-121, to generate two global feature spaces. For the stylized CXR, another round of style randomization is applied to the feature space after dense block-2, with a feature space for a randomly selected stylized CXR from the training mini-batch. The content consistency loss is employed between the two global feature spaces to increase the model’s bias toward disease-specific content. In addition, a Kullback-Leibler divergence-based regularization loss is applied between the predicted probability distributions for CXRs with and without style statistics perturbed. The global feature space from the stylized CXR is pooled and passed to the classifier for pathology prediction.
@@ III-B3 Kullback-Leibler Divergence loss
The Kullbeck-Leibler divergence (KLD) loss is generally applied to minimize the difference between two distributions [45]. For example, if we try to reduce the distribution gap between distributions, $\mathbf{y_{pred}}$ and $\mathbf{y_{true}}$, the loss is defined as:
$\operatorname{KLD}(\mathbf{y_{pred}},\mathbf{y_{true}})$ $=\mathbf{y_{true}}\log{\frac{\mathbf{y_{true}}}{\mathbf{y_{pred}}}}$ (5)
Where $\mathbf{y_{true}}$ is the target distribution and $\mathbf{y_{pred}}$ is generally the output from a model. In our proposed method, we will derive two predictive distributions, $\mathbf{y_{1}}$, and $\mathbf{y_{2}}$, one for images with style perturbed and the other without style perturbed. As there are no ground truth predictive distributions, to minimize the discrepancy between these predicted predictive distributions, we define the loss as:
$\operatorname{KLD}^{\prime}(\mathbf{y_{1}},\mathbf{y_{2}})$ $=\frac{\mathbf{y_{1}}^{*}\log{\frac{\mathbf{y_{1}}^{*}}{\mathbf{y_{2}}}}+\mathbf{y_{2}}^{*}\log{\frac{\mathbf{y_{2}}^{*}}{\mathbf{y_{1}}}}}{2}$ (6)
@@ III-D Feature-Level Style Perturbation (SRM-FL)
In medical image segmentation, it has been demonstrated that multiple stacked augmentations or strong augmentations can effectively improve a model’s robustness to domain shift [13]. Inspired by this finding, we apply the style perturbation not only at the image level but also at the feature level. The feature-level style randomization module takes content ($\mathbf{X_{c}}\in\mathbb{R}^{D\times W\times H}$) and reference style ($\mathbf{X_{rs}}\in\mathbb{R}^{D\times W\times H}$) activation maps and transfers the style of the content activation maps to that of the style activation maps. We deploy two style learning nets, i.e., GammaNet ($\phi_{\gamma}$) and BetaNet ($\phi_{\beta}$), to learn the pixel-level affine transformation parameters for transferring style to the content activation maps. The GammaNet and BetaNet shares similar architecture, comprised of four convolutional layers. The initial and the last convolutional layers contain kernels of 1×1 with the same number of channels as the input feature map. The second and third convolutional layers have filters with 3×3 dimensions, with channel dimensions following the squeeze and excitation block structure [47], i.e., the second block reduces the channel number, while the third block expands the channel number. The style nets do not contain any non-linear activation functions or normalization layers [48]. The output from these two nets, $\phi_{\gamma}(\mathbf{X_{rs}})\in\mathbb{R}^{D\times W\times H}$, and $\phi_{\beta}(\mathbf{X_{rs}})\in\mathbb{R}^{D\times W\times H}$, are set as the affine transformation parameters in the IN function to transfer the style statistics.
@@ III-E Domain Agnostic Network
The overall proposed framework is illustrated in Fig. 2. We adopt DenseIBN-121 [50], pre-trained on the ImageNet dataset, as the backbone of our framework. Furthermore, we replace the last fully connected (FC) layer with an FC layer with $N$ output neurons, where $N$ denotes the class number.
@@
$\mathcal{L}_{pdr}=\operatorname{KLD}^{\prime}(\mathbf{p^{s}},\mathbf{p})$ (17)
@@
The overall learning objective for the backbone model (without style nets) is comprised of the classification loss and consistency regularization loss and is expressed as:
@@
Training configs Values
Weight initialization pre-trained on ImageNet
Optimizer Adam
Optimizer momentum $\beta_{1}$=0.9, $\beta_{2}$=0.999
Learning rate 0.0001
Batch size 50
Gradient accumulation step 4
Training epochs 50
Input image dimension 224×224
Augmentations Horizontal flip and Random cropping [1]
Classification Loss Focal loss with $\alpha^{t}$=0.25 and $\gamma^{\prime}$=2 [51]
Exp. Mov. Avg. (EMA) [52] 0.997
@@ IV-A Thoracic Disease Datasets
We utilize the five large-scale thoracic disease datasets, namely, CheXpert [23], MIMIC-CXR [24], BRAX [25], NIH chest X-ray14 [31], and VinDr-CXR [32] datasets, for our experiments. The CheXpert, MIMIC-CXR, and BRAX datasets obtain 224,316 CXRs from 65,240 patients, 377,110 CXRs from 227,835 imaging studies for 65,379 patients, and 40,967 CXRs from 24,959 clinical studies, respectively. These three datasets have the same label space consisting of 14 pathologies. We follow the U-Ones policy [23] to prepare the ground truths. We use frontal CXR images and select one image per patient/study for our experiments. The VinDr-CXR dataset contains 18,000 images manually annotated by a total of 17 experienced radiologists. The NIH chest X-ray14 dataset comprises 112,120 chest X-rays from 30,805 unique patients with 14 diseases. The VinDr-CXR and NIH chest X-ray14 datasets have 10 and 7 common labels, respectively, with the CheXpert, MIMIC-CXR, and BRAX datasets. We have utilized these common labels for reporting the performance of these two datasets [12]. We use CheXpert and MIMIC-CXR datasets for training and the BRAX, NIH chest X-ray14, and VinDr-CXR datasets for unseen domain inference.
@@ IV-B Training Scheme
We initialize the backbone weights with DenseIBN-121 [50] pre-trained on ImageNet. We resize the CXR images to 256×256 and normalize them with the mean and standard deviation of the ImageNet training set. We randomly crop 224×224 patches during training and use a centrally cropped sub-image of 224×224 dimensions for validation and inference. We provide the summary of training hyperparameter settings in Table II. We progressively add each proposed component to the base model and utilize the SRM modules only during training time. We employ the percentage area under the receiver operating characteristic curve (AUC) for performance evaluation [9, 12]. For comparison with other methods, we utilize the multi-label stratified k-fold scheme [53] [^1] https://github.com/trent-b/iterative-stratification to split the training dataset into a 5-fold cross-validation framework to report the mean and standard deviation of the evaluation metrics across the folds. We use the paired t-test between cross-validated mean AUCs for reporting statistical significance.
@@ IV-C2 Single-source Domain Generalization
TABLE III: Performance Comparison of the Proposed Method with State-of-the-art Systems on the Unseen Domain Test Datasets, i.e., BRAX, NIH Chest X-ray14, and VinDr-CXR, Under the Multi-source Domain Generalization Setting. The Best Result is Shown in Red.
TABLE IV: Performance Comparison of the Proposed Method with State-of-the-art Systems on the Unseen Domain Test Datasets, i.e., BRAX, NIH Chest X-ray14, and VinDr-CXR, Under the Single-source Domain Generalization Setting. The Best Result is Shown in Red.
@@ IV-D Analysis of Different Components
We conduct experiments to evaluate the effect of each proposed component on the unseen domain test datasets, including image-level style perturbation (SRM-IL), feature-level style perturbation (SRM-FL), and consistency regularization loss ($\mathcal{L}_{cons}$). The results are reported in Table V. Checkmarks in the table indicate if the specific component is incorporated. As shown in Table V, although all three components improve the model performance, SRM-IL and consistency regularization loss contribute more strongly. This stronger contribution is expected since consistency loss reinforces the feature extractor to focus on the content region, and the SRM-IL module increases diversity.
We have conducted an ablation study to evaluate the impact of content consistency regularization ($\mathbf{\mathcal{L}_{ccr}}$) and predictive distribution regularization ($\mathbf{\mathcal{L}_{pdr}}$) losses on classification performance. The results are presented in Table VI. Combining both of the consistency regularization losses results in 77.32±0.35, 82.63±0.13, and 88.38±0.19 average AUC (%) score on the unseen domain test datasets, i.e., BRAX, NIH Chest X-ray14, and VinDr-CXR datasets. These scores surpass the performance level achieved by using either of the regularization loss alone. We hypothesize that while predictive distribution regularization loss enforces consistency on two probability distributions from two different style augmented versions of the same CXR image, it does not guarantee that the backbone model will focus on the same spatial region for predicting the consistent probabilities. In order to address this issue, we employ the content consistency loss in addition to the predictive distribution loss. We can also observe that the content consistency loss results in better performance compared to the predictive distribution loss when they are utilized alone.
@@ IV-E Analysis of SRM-FL
Indices Base SRM-IL SRM-FL $\mathcal{L}_{cons}$ BRAX NIH Chest X-Ray14 VinDr-CXR (i) 73.08±0.75 81.12±0.13 84.75±0.27 (ii) 76.13±0.62 81.75±0.15 86.30±0.30 (iii) 76.24±0.51 81.91±0.09 86.43±0.59 (iv) 75.19±0.89 81.51±0.26 86.25±0.40 (v) 77.20±0.36 82.41±0.06 87.68±0.30 (vi) 77.32±0.35 82.63±0.13 88.38±0.19
TABLE VI: Ablation Study on the Consistency Regularization Losses. The Best Result is Shown in Red.
Indices $\mathcal{L}_{ccr}$ $\mathcal{L}_{pdr}$ BRAX NIH Chest X-ray14 VinDr-CXR (i) 76.24±0.51 81.91±0.09 86.43±0.59 (ii) 76.68±0.44 82.22±0.08 87.43±0.58 (iii) 76.93±0.31 82.31±0.04 88.37±0.34 (iv) 77.32±0.35 82.63±0.13 88.38±0.19
TABLE VII: Analysis of the Feature-Level Style Randomization Module. The Best Result for Each Disease Class is Shown in Red.
Part-1: Investigation of the effect of randomization stage
Method BRAX NIH Chest X-Ray14 VinDr-CXR
DB-1 77.00±0.39 82.34±0.08 87.63±0.63
DB-2 77.32±0.35 82.63±0.13 88.38±0.19
DB-3 76.75±0.44 82.34±0.13 88.15±0.35
Part-2: Effect of learnable style embeddings
Method BRAX NIH Chest X-Ray14 VinDr-CXR
Pre-defined 77.04±0.48 82.44±0.13 88.09±0.48
Learnable 77.32±0.35 82.63±0.13 88.38±0.19
Part-3: Analysis of the style net architecture
Method BRAX NIH Chest X-Ray14 VinDr-CXR
Conv 77.32±0.35 82.63±0.13 88.38±0.19
Conv+ReLU 77.17±0.30 82.30±0.10 88.23±0.30
Conv+BN 77.26±0.40 82.38±0.14 88.19±0.23
TABLE VIII: Investigation of the Impact of Different Input CXR Dimensions on the BRAX ${}^{\dagger}$. The Best Result for Each Disease Class is Shown in Red.
@@
Sensitivity of the hyperparameter $\boldsymbol{\eta}$: We vary the hyper-parameter $\eta$ from 1 to 1e-5 and calculate the percentage AUC performance. The results are demonstrated in Fig. 3. From Fig. 3, we can observe that increasing $\eta$ too much degrades the score as it disrupts the semantic content features. Decreasing the $\eta$ too much reduces the style diversity of the SRM-FL, resulting in performance loss. The hyper-parameter $\eta$ controls the stylization capacity of the style nets by governing the content and style loss. CXRs often have complex backgrounds, and too-hard style perturbations may increase the complexity and potential interference with diagnosing small lesions. To verify this, we plot AUC (%) scores of two different lesions, i.e., cardiomegaly and lung lesions, compared to the $\eta$ values, on the unseen domain BRAX test dataset. Note that the small lesions, i.e., masses and nodules, are included in the lung lesion category. The plots are given in Fig. 4. We can observe that for large lesions such as cardiomegaly, the AUC (%) does not vary much concerning the $\eta$ value. However, the $\eta$ value significantly impacts the detection of the lung lesion category, i.e., small lesions. Therefore, it is recommended to calculate the optimal $\eta$ value for performing the feature-level style perturbation. Another strategy that can be adopted is to increase the spatial dimension of the input image, which is often utilized in literature for thoracic disease detection to improve the detection of small lesions [1]. To demonstrate this, we have trained our model on 384×384 input images and compared the performance with the 224×224 input images under $\eta$=0.01, on the unseen domain BRAX test dataset. The results are reported in Table VIII. We can observe that the AUC (%) of the lung lesion class improves from 60.44 to 65.32.

### 25 Feb 2024 v4

@@ I Introduction
In this work, we propose a novel domain-agnostic, content-biased, and style-invariant DG framework that can be trained in single or multi-source domain settings without explicit domain labels. We employ style randomization modules (SRMs) at the image and feature levels to hierarchically augment the style while preserving the pathology-specific content traits. Utilizing the prior maximum and minimum pixel value knowledge of a normalized CXR image, we construct a set for sampling mean and standard deviation to use them for switching style characteristics at the image level. For feature-level style augmentation, we deploy a learnable SRM module that generates the pixel-wise style embeddings instead of setting pre-defined parameters, i.e., channel-wise mean and standard deviations as affine parameter values [27], for style manipulation. In addition, we employ two consistency regularization losses to enforce the consistency of semantic global feature maps and the predicted probability distributions, respectively, between the original CXR and the style-perturbed CXR. The variability of styles between different perturbations of the same CXR with content information intact makes the style attribute unreliable and makes the model focus more on pathology-specific content in making a decision. We also share the source code [^1] https://github.com/mHealthBuet/DomainAgnosticCXR of our method for further dissemination of our approach and experiments. The contributions of this paper are summarized as follows:
@@ II Related work
The feature-level randomization includes the perturbation of feature statistics from intermediate convolutional layers of the backbone model. Nuriel et al.[29] swapped the style statistics between activations of permuted and non-permuted samples in a mini-batch to reduce the representation of global statistics in the classifier model. However, they merely swap the style statistics, limiting the sample diversity to the available source domains. Zhou et al.[19] proposed a novel parameter-free module MixStyle that synthesizes new domains by probabilistically mixing feature statistics between instances. Nevertheless, insufficient diversity still exists due to reliance on source domains. Nam et al.[21] utilized adaptive instance normalization along with adversarial learning to reduce the intrinsic style bias. However, the style-biased network increases training cost and time. Tang et al.[30] proposed two novel modules named CrossNorm and SelfNorm for swapping feature statistics and recalibrating the statistics through attention to both diversify and bridge the gap between source and target domains. Though the SelfNorm is dynamic, the CrossNorm is limited to training data. Wang et al.[22] developed an encoder-decoder-like architecture named FSR to integrate random noise with the original style parameters to generate new style embeddings. However, they use the pre-defined parameters, i.e., channel-wise mean and standard deviation, as the style embeddings. Instead of pre-defined parameters, pixel-wise learnable style parameters may yield more diverse and hard style statistics, resulting in a more robust DG model.
@@ III-B1 Instance Normalization
Figure 2: Overview of the proposed framework. The style statistics of the input CXR are randomized with the randomly sampled mean and standard deviation from the set of values constructed from the prior knowledge of possible maximum and minimum values. Both CXRs, with or without style randomized, are passed to the shared feature extractor, DenseIBN-121, to generate two global feature spaces. For the stylized CXR, another round of style randomization is applied to the feature space after dense block-2, with a feature space for a randomly selected stylized CXR from the training mini-batch. The feature-level style randomization block has learnable parameters and is trained alongside the backbone model. While the image-level style randomization block has no learnable parameters, it achieves diversity by sampling different style statistics per CXR image in each iteration for every epoch. The content consistency loss is employed between the two global feature spaces to increase the model’s bias toward disease-specific content. In addition, a Kullback-Leibler divergence-based regularization loss is applied between the predicted probability distributions for CXRs with and without style statistics perturbed. The global feature space from the stylized CXR is pooled and passed to the classifier for pathology prediction.
@@ III-B3 Kullback-Leibler Divergence loss
The Kullbeck-Leibler divergence ($D_{\text{KL}}(\mathbf{P}||\mathbf{Q})$) loss is applied to minimize the difference between two probability distributions [45], where $\mathbf{P}$ is the target, and $\mathbf{Q}$ is generally the output from a model. In our proposed method, we will derive two predictive distributions, $\mathbf{y_{1}}$, and $\mathbf{y_{2}}$, one for images with style perturbed and the other without style perturbed. As there are no ground truth predictive distributions, to minimize the discrepancy between these predicted predictive distributions, we define the loss as:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}D_{\text{KL}}^{\prime}(\mathbf{y_{1}},\mathbf{y_{2}})$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}=\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\frac{D_{\text{KL}}(\mathbf{y_{1}}^{*}||\mathbf{y_{2}})+D_{\text{KL}}(\mathbf{y_{2}}^{*}||\mathbf{y_{1}})}{2}$ (5)
@@ III-D Feature-Level Style Perturbation (SRM-FL)
In medical image segmentation, it has been demonstrated that multiple stacked augmentations or strong augmentations can effectively improve a model’s robustness to domain shift [13]. Inspired by this finding, we apply the style perturbation not only at the image level but also at the feature level. The feature-level style randomization module takes content ($\mathbf{X_{c}}\in\mathbb{R}^{D\times W\times H}$) and reference style ($\mathbf{X_{rs}}\in\mathbb{R}^{D\times W\times H}$) activation maps and transfers the style of the content activation maps to that of the style activation maps. We deploy two style learning nets, i.e., GammaNet ($\phi_{\gamma}$) and BetaNet ($\phi_{\beta}$), to learn the pixel-level affine transformation parameters for transferring style to the content activation maps. The GammaNet and BetaNet shares similar architecture, comprised of four convolutional layers. The initial and the last convolutional layers contain kernels of 1×1 with the same number of channels as the input feature map. The second and third convolutional layers have filters with 3×3 dimensions, with channel dimensions following the squeeze and excitation block structure [47], i.e., the second block reduces the channel number, while the third block expands the channel number. The style nets do not contain any non-linear activation functions or normalization layers. The output from these two nets, $\phi_{\gamma}(\mathbf{X_{rs}})\in\mathbb{R}^{D\times W\times H}$, and $\phi_{\beta}(\mathbf{X_{rs}})\in\mathbb{R}^{D\times W\times H}$, are set as the affine transformation parameters in the IN function to transfer the style statistics.
@@ III-E Domain Agnostic Network
The overall proposed framework is illustrated in Fig. 2. We adopt DenseIBN-121 [49] as the backbone of our framework [9]. Furthermore, we replace the last fully connected (FC) layer with an FC layer with $N$ output neurons, where $N$ denotes the class number. Please note that our proposed approach is architecture-agnostic and can also be extended to other CNN/Transformer-based backbones. Please refer to subsection A of the supplementary material for experimental results on different backbone architectures.
@@
$\mathcal{L}_{pdr}=D_{\text{KL}}^{\prime}(\mathbf{p^{s}},\mathbf{p})$ (16)
@@
The overall learning objective for the backbone model (without style nets) is expressed as:
@@
Training configs Values Weight initialization pre-trained on ImageNet Optimizer Adam Optimizer momentum $\beta_{1}$=0.9, $\beta_{2}$=0.999 Learning rate 0.0001 Batch size 50 Gradient accumulation step 4 Training epochs 50 Input image dimension 224×224 Augmentations Horizontal flip and Random cropping [1] Classification Loss Focal loss with $\alpha^{t}$=0.25 and $\gamma^{\prime}$=2 [50] Exp. Mov. Avg. (EMA) [51] 0.997
TABLE III: Performance Comparison with State-of-the-art Systems on the Unseen Domain Test Datasets, i.e., BRAX, NIH Chest X-ray14, and VinDr-CXR, Under the Multi-source Domain Generalization Setting. The Best Result is Shown in Red.
@@ IV-A Thoracic Disease Datasets
We utilize the five large-scale thoracic disease datasets, namely, CheXpert [23], MIMIC-CXR [24], BRAX [25], NIH chest X-ray14 [31], and VinDr-CXR [32] datasets, for our experiments. The CheXpert, MIMIC-CXR, and BRAX datasets have the same label space consisting of 14 pathologies. We follow the U-Ones policy [23] to prepare the ground truths. We use frontal CXR images and select one image per patient/study for our experiments. The VinDr-CXR and NIH chest X-ray14 datasets have 10 and 7 common labels, respectively, with the CheXpert, MIMIC-CXR, and BRAX datasets. We have utilized these common labels for reporting the performance of these two datasets [12]. We use CheXpert and MIMIC-CXR datasets for training and the BRAX, NIH chest X-ray14, and VinDr-CXR datasets for unseen domain inference.
@@ IV-B Training Scheme
We resize the CXR images to 256×256 and normalize them with the mean and standard deviation of the ImageNet training set. We randomly crop 224×224 patches during training and use a centrally cropped sub-image of 224×224 dimensions for validation and inference. We provide the summary of training hyperparameter settings in Table II. We progressively add each proposed component to the base model and utilize the SRM modules only during training time. We employ the percentage area under the receiver operating characteristic curve (AUC) for performance evaluation [9, 12]. For comparison with other methods, we utilize the multi-label stratified k-fold scheme [52] to split the training dataset into a 5-fold cross-validation framework to report the mean and standard deviation of the evaluation metrics across the folds. We use the paired t-test between cross-validated mean AUCs for reporting statistical significance.
@@ IV-C1 Multi-source Domain Generalization
TABLE IV: Performance Comparison with State-of-the-art Systems on the Unseen Domain Test Datasets, i.e., BRAX, NIH Chest X-ray14, and VinDr-CXR, Under the Single-source Domain Generalization Setting. The Best Result is Shown in Red.
@@ IV-D Analysis of Different Components
We conduct experiments to evaluate the effect of each proposed component on the unseen domain test datasets, including image-level style perturbation (SRM-IL), feature-level style perturbation (SRM-FL), and consistency regularization losses ($\mathcal{L}_{ccr}$ and $\mathcal{L}_{pdr}$). The results are reported in Table V. Checkmarks in the table indicate if the specific component is incorporated. As shown in Table V, although all components improve the model performance, SRM-IL and consistency regularization losses contribute more strongly. This stronger contribution is expected since consistency losses reinforce the feature extractor to focus on the content region, and the SRM-IL module increases diversity.
We can also observe that the scores achieved by the combination of content consistency regularization ($\mathcal{L}_{ccr}$) and predictive distribution regularization ($\mathcal{L}_{pdr}$) losses surpass the performance level achieved by using either of the regularization loss alone (indices-viii, ix, x). We hypothesize that while predictive distribution regularization loss enforces consistency on two probability distributions from two different style augmented versions of the same CXR image, it does not guarantee that the backbone model will focus on the same spatial region for predicting the consistent probabilities. In order to address this issue, we employ the content consistency loss in addition to the predictive distribution loss. We can also observe that the content consistency loss results in better performance compared to the predictive distribution loss when they are utilized alone.
We also conduct an ablation study on the symmetric design of the predictive distribution regularization loss. The results are reported in Table VI. We can observe that utilizing the symmetric design achieves the best scores (indices-iv) while optimizing only one part does not ensure optimal performance (indices-ii and iii).
@@ IV-E Analysis of SRM-FL
Indices Base SRM-IL SRM-FL $\mathcal{L}_{ccr}$ $\mathcal{L}_{pdr}$ BRAX NIH Chest X-Ray14 VinDr-CXR (i) 73.08±0.75 81.12±0.13 84.75±0.27 (ii) 76.13±0.62 81.91±0.09 86.43±0.59 (iii) 76.24±0.51 82.03±0.06 86.56±0.65 (iv) 75.19±0.89 81.51±0.26 86.25±0.40 (v) 77.06±0.29 82.46±0.12 87.96±0.46 (vi) 76.38±0.40 82.30±0.11 86.67±0.73 (vii) 77.20±0.36 82.47±0.06 87.68±0.30 (viii) 76.93±0.31 82.31±0.04 88.37±0.34 (ix) 76.68±0.44 82.22±0.08 87.43±0.58 (x) 77.32±0.35 82.63±0.13 88.38±0.19
TABLE VI: Ablation Study on the Symmetric Design of the $\mathcal{L}_{pdr}$ Loss. The Best Result is Shown in Red.
Indices Optimize y2 Optimize y1 BRAX NIH Chest X-ray14 VinDr-CXR (i) 76.93±0.31 82.31±0.04 88.37±0.34 (ii) 77.00±0.19 82.24±0.10 87.64±0.44 (iii) 76.86±0.24 82.32±0.10 87.90±0.31 (iv) 77.32±0.35 82.63±0.13 88.38±0.19
TABLE VII: Analysis of the Feature-Level Style Randomization Module. The Best Result is Shown in Red.
Part-1: Investigation of the effect of randomization stage Method BRAX NIH Chest X-Ray14 VinDr-CXR DB-1 77.00±0.39 82.34±0.08 87.63±0.63 DB-2 77.32±0.35 82.63±0.13 88.38±0.19 DB-3 76.75±0.44 82.34±0.13 88.15±0.35 Part-2: Effect of learnable style embeddings Method BRAX NIH Chest X-Ray14 VinDr-CXR Pre-defined 77.04±0.48 82.44±0.13 88.09±0.48 Learnable 77.32±0.35 82.63±0.13 88.38±0.19 Part-3: Analysis of the style net architecture Method BRAX NIH Chest X-Ray14 VinDr-CXR Conv 77.32±0.35 82.63±0.13 88.38±0.19 Conv+ReLU 77.17±0.30 82.30±0.10 88.23±0.30 Conv+BN 77.26±0.40 82.38±0.14 88.19±0.23 Part-4: Effect of style source selection procedure Method BRAX NIH Chest X-Ray14 VinDr-CXR Class-balanced Sampling 77.15±0.29 82.39±0.12 88.18±0.46 Permuted-mini Batch 77.32±0.35 82.63±0.13 88.38±0.19
TABLE VIII: Investigation of the Impact of Different Input CXR Dimensions on the BRAX ${}^{\dagger}$. The Best Result is Shown in Red.
@@
Effect of style source selection procedure: To evaluate the impact of the style source selection procedure, we construct the style source for each iteration from the training batch by balance sampling of all 14 classes and compare performance with the permuted mini-batch style source selection process. The results are reported in part-4 of Table VII. We can observe that the permuted-mini batch as style source achieves the best score compared to the class-balanced sampling of style source, which can be attributed to the fact that the oversampling of classes with fewer examples results in the model being exposed to fewer style-diverse samples.
Sensitivity of the hyperparameter $\boldsymbol{\eta}$: We vary the hyper-parameter $\eta$ from 1 to 1e-5 and calculate the percentage AUC performance. The results are demonstrated in Fig. 3. From Fig. 3, we can observe that increasing $\eta$ too much degrades the score as it disrupts the semantic content features. Decreasing the $\eta$ too much reduces the style diversity of the SRM-FL, resulting in performance loss. The hyper-parameter $\eta$ controls the stylization capacity of the style nets by governing the content and style loss. CXRs often have complex backgrounds, and too-hard style perturbations may increase the complexity and potential interference with diagnosing small lesions. To verify this, we plot AUC (%) scores of two different lesions, i.e., cardiomegaly and lung lesions, compared to the $\eta$ values, on the unseen domain BRAX test dataset. Note that the small lesions, i.e., masses and nodules, are included in the lung lesion category. The plots are given in Fig. 4. We can observe that for large lesions such as cardiomegaly, the AUC (%) does not vary much concerning the $\eta$ value. However, the $\eta$ value significantly impacts the detection of the lung lesion category, i.e., small lesions. Therefore, it is recommended to calculate the optimal $\eta$ value for performing the feature-level style perturbation. Another strategy that can be adopted is to increase the spatial dimension of the input image, which is often utilized in literature for thoracic disease detection to improve the detection of small lesions [1]. To demonstrate this, we have trained our model on 384×384 input images and compared the performance with the 224×224 input images under $\eta$=0.01, on the unseen domain BRAX test dataset. The results are reported in Table VIII. We can observe that the AUC (%) of the lung lesion improves from 60.44 to 65.32.